OpenStudy (anonymous):

(x+2)/7 - (3x-1)/(4x) Simplify rational expression

jhonyy9 (jhonyy9):

(4x(x+2))-(7(3x-1)) /28x =(4x*2+8x-21x+7)/28x =(4x*2-13x+7)/28x

OpenStudy (anonymous):

=x/7-13/28+1/4x =(4x^2-13x+7)/28x